1.     FET’s preferred stock is selling at $44 on the market and pays an annual dividend of $3.20 per share.

a.     What is the expected rate of return on the stock?

b.     If an investor's required rate of return is 10%, what is the value of the stock to that investor?

c.     Considering the investor's required rate of return, does this stock seem to be a desirable investment?

Please share the details...


1
Expert's answer
2021-04-22T06:15:45-0400

Part A:

A Rate of Return (ROR) is the increase or loss of a venture over a definite time. The rate of yield is the increase (or loss) associated with the budget of an initial venture, naturally articulated in the form of a fraction. When the ROR is positive, it is reflected as an increase, and when the ROR is negative, it redirects a loss on the venture.

Bear in mind that the business must involve any increase made during the holding period of the venture within the formula; in this question, no data regarding the value of the stock. Thus, we have to include the dividend merely to calculate.

ROR = 3.2/ 44 = 7.27 %.




Part B:

Value of stock = Dividend or Coupon / R

32 = 3.2 / 10%

Part C:

The actual stockholder rate of return; this not an appropriate venture as the stakeholder requires a 10 % return, but the projected return is simply 7. 27 %.